OpenStudy (anonymous):

Math Review, I got alot of question on my test and homework this week, need someone to solve these for me so i know what i did wrong please

OpenStudy (anonymous):

Find the Slope m= y2 - y1/ X2 - X1 of a line passing through the given points (-4,2) and (1,-3) Find the Slope m= y2 - y1/ X2 - X1 of a line passing through the given points (3,3) and (4,0)

OpenStudy (anonymous):

where is your work?

OpenStudy (anonymous):

Hey, you ask for checking not doing. I need your work

OpenStudy (anonymous):

i dnt have it srry

OpenStudy (anonymous):

this is just a paper that my teacher gave me today to review

OpenStudy (anonymous):

Hope this helps :)

OpenStudy (anonymous):

these question were on my test that i took on wednesday

OpenStudy (anonymous):

ok, Fizzy is right

OpenStudy (anonymous):

thnks but can you explain how to do it @FizzyC

OpenStudy (anonymous):

i need help with these as well What is the slope of the line represented by the equation y = 7? What is the slope of the line represented by the equation x = -9? Write an equation of a line with a slope of 0 and passing through the point (5,4). Write an equation of a line with an undefined slope and passing through the point (-2,4). Write an equation of a line that passes through (9,-5) and (3,-5). Write an equation of a line that passes through the points (2,3) and (2,-6).

OpenStudy (anonymous):

Sure :) well the equation to find the gradient given two coordinates is just like you said: \[m=\frac{y _{2} -y _{1} }{x _{2} -x _{1} }\] so you have two coordinates, (-4;2) & (1;-3) and you know they in the form (x;y), so take any coordinate to start with. y2-y1 means take the y-value from one coordiante and subtract it from the y-value of the other coordinate. so on the top of the fraction we take 2-(-3) now note the order did not matter I could have said -3-2, ONLY as long as you do the same order for the bottom. the order of which coordinate you start with does not matter, as long as the order you subtract the y terms are consistent with the x terms, you got no problems. Hope this makes sense? Now for your next questions: What is the slope of the line represented by the equation y = 7? --> Well the equation for a line is y=mx+c where m is gradient and c the value at which the line cuts the y axis. so in your question you have something like this y=0x +7, where your gradient is zero and hence 0*x is just zero, you are just left with y =7. I'll add a graphic to demonstrate. What is the slope of the line represented by the equation x = -9? --> Well x=-9 is a vertical line passing through the x value of -9, we refer to this gradient as undefined, i'll add sketch as well. Write an equation of a line with a slope of 0 and passing through the point (5,4). --> well y=mx+c is the general equation, so m =slope, and in your case slope is 0, so m=0, so we get y=0(x)+c, now find c by substituting the given point (5,4), therefore you get 4=0(5)+C and so c=4, note how i just kept the zero gradient multiplied by the x, just to show clearly how you would do it generally. so we get y=0x+4 which is just y=4 Write an equation of a line with an undefined slope and passing through the point (-2,4). --> very similar to the one of the previous questions, and think about this after what I explained in that question, when the slope is undefined you have a straight vertical line passing through the x coordinate given, so in this case x=-2 is your equation. note that the y-value of 4 is not really important here. look at drawing 3 in my attached file to see why Write an equation of a line that passes through (9,-5) and (3,-5). --> y=mx+c find gradient, m=(y2-y1)/(x2-x1) so\[m=\frac{ -5-(-5) }{ 9-3 }\] which is: \[\frac{ 0 }{ 6 }\] and that is just zero, so m=0, so y=0(x)+c, now substitute any point to find C, i'm taking (9,-5) so you get: -5=0(9)+C so C=-5 therefore y=0(x)-5 which is just y=-5 (a straight horizontal line) Write an equation of a line that passes through the points (2,3) and (2,-6). --> same method as previous Hope this helps :) and makes sense.

OpenStudy (anonymous):

i get it kinda thanks im gonna try your method on some of my homework problems

OpenStudy (anonymous):

ok for the first questions i asked how do you get M= -3 -2/ 1-(-4) = -5/5 =-1

OpenStudy (anonymous):

I just plug the x and y values of the two coordinates into that formula. check out this video, it explains very well. there are more topics on his website too, so you can browse through them, maybe they help. https://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/slope-and-intercepts/v/slope-of-a-line-2

OpenStudy (anonymous):

thnks man now i see how to do it

OpenStudy (anonymous):

ok for Write an equation of a line that passes through (9,-5) and (3,-5). where do you get y=mx+c from